OpenStudy (anonymous):

rtex of the quadratic equation: y = 2x2 + 12x − 4.

OpenStudy (anonymous):

rtex?

OpenStudy (anonymous):

vertex

OpenStudy (anonymous):

2(x^2 + 6x-2)=0 x^2+6x-2=0 x = (-6+(36+8)^(1/2) / 2 x= some complex

OpenStudy (anonymous):

what is a vertex ?

OpenStudy (anonymous):

yes i need coordinates

OpenStudy (anonymous):

let the vertex=(x,y)

OpenStudy (anonymous):

then x=-b/za, according to y=ax^2+bx+c

OpenStudy (anonymous):

x=-b/2a

OpenStudy (anonymous):

so the answer is (-3,-22), do you understand?

OpenStudy (anonymous):

f'(x)=2(2x+6)=0 x = -3 y = -22 f''(x)=4 >0 hence local minimum (-3,-22) ??

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

yes ,the minimum is -22